OpenStudy (dls):

How many symmetric matrices can be obtained by using 0,0,0,0,0,-1,-1,1,1

OpenStudy (anonymous):

3

OpenStudy (anonymous):

ooh nope, it's more

OpenStudy (dls):

yup,its not 3

OpenStudy (anonymous):

6, the 1's have to be on a diagonal

OpenStudy (dls):

I don't have answer for this one,I have a similar question with the correct answer so I am changing the values,tell now

OpenStudy (anonymous):

ooops. 6 isn't enough either...

OpenStudy (dls):

stop guessing :|

OpenStudy (anonymous):

okay, both 2's or none can be on the diagonal. likewise, two 0's or none on the diagonal. so one 1 or all three 1's must be on a diagonal. right?

OpenStudy (dls):

I changed the values! there are no 2's now,this is easier.

OpenStudy (anonymous):

2(6+9+9) = 48

OpenStudy (dls):

wrong again

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

sorry, duplicates when 0 is in the 2,2 position

OpenStudy (dls):

Its 12.

OpenStudy (anonymous):

\[\left[\begin{matrix}0 & 0 & -1 \\0 & 0 & 1\\ -1 &1 & 0\end{matrix}\right] \left[\begin{matrix}0 & 0 & 1 \\0 & 0 & =1\\ 1 &=1 & 0\end{matrix}\right]\] \[\left[\begin{matrix}0 & 1 & -1 \\1 & 0 & 0\\ -1 &0 & 0\end{matrix}\right] \left[\begin{matrix}0 & -1 & 1 \\-1 & 0 & 0\\ 1 &0 & 0\end{matrix}\right]\] \[\left[\begin{matrix}0 & 1 & 0 \\1 & 0 & -1\\ 0 &-1 & 0\end{matrix}\right] \left[\begin{matrix}0 & -1 & 0 \\-1 & 0 & 1\\ 0 &1 & 0\end{matrix}\right]\] \[\left[\begin{matrix}0 & 0 & 0 \\0 & 1 & -1\\ 0 &-1 & 1\end{matrix}\right] \left[\begin{matrix}0 & -1 & 0 \\-1 & 1 & 0\\ 0 &0 & 1\end{matrix}\right]\left[\begin{matrix}0 & 0 & -1 \\0 & 1 & 0\\ -1 &0 & 1\end{matrix}\right]\] \[\left[\begin{matrix}1 & 0 & 0 \\0 & 0 & -1\\ 0 &-1 & 1\end{matrix}\right] \left[\begin{matrix}1 & -1 & 0 \\-1 & 0 & 0\\ 0 &0 & 1\end{matrix}\right]\left[\begin{matrix}1 & 0 & -1 \\0 & 0 & 0\\ -1 &0 & 1\end{matrix}\right]\] \[\left[\begin{matrix}1 & 0 & 0 \\0 & 1 & -1\\ 0 &-1 & 0\end{matrix}\right] \left[\begin{matrix}1 & -1 & 0 \\-1 & 1 & 0\\ 0 &0 & 0\end{matrix}\right]\left[\begin{matrix}1 & 0 & -1 \\0 & 1 & 0\\ -1 &0 & 0\end{matrix}\right]\] \[\left[\begin{matrix}0 & 0 & 0 \\0 & -1 & 1\\ 0 & 1 & -1\end{matrix}\right] \left[\begin{matrix}0 & 1 & 0 \\1 & -1 & 0\\ 0 &0 & -1\end{matrix}\right]\left[\begin{matrix}0 & 0 & 1 \\0 & -1 & 0\\ 1 &0 & -1\end{matrix}\right]\] \[\left[\begin{matrix}-1 & 0 & 0 \\0 & 0 & 1\\ 0 &1 & -1\end{matrix}\right] \left[\begin{matrix}-1 & 1 & 0 \\1 & 0 & 0\\ 0 &0 & -1\end{matrix}\right]\left[\begin{matrix}-1 & 0 & 1 \\0 & 0 & 0\\ 1 &0 & -1\end{matrix}\right]\] \[\left[\begin{matrix}-1 & 0 & 0 \\0 & -1 & 1\\ 0 &1 & 0\end{matrix}\right] \left[\begin{matrix}-1 & 1 & 0 \\1 & -1 & 0\\ 0 &0 & 0\end{matrix}\right]\left[\begin{matrix}-1 & 0 & 1 \\0 & -1 & 0\\ 1 &0 & 0\end{matrix}\right]\] ummm. it's more than 12

OpenStudy (anonymous):

Assign the 1s, then assign the -1s... the zeroes will fill the remaining parts in.

OpenStudy (anonymous):

Also, since it is symmetric, only assign the bottom left of the triangle..

OpenStudy (dls):

Yeah,@wio the solution is something like that,but I need a little detailed explanation..just one example of how we do it,like for one matrix..

OpenStudy (dls):

@terenzreignz @oldrin.bataku

OpenStudy (anonymous):

Do you know what a symmetric matrix is?

OpenStudy (dls):

\[\Huge a_{ij}=a_{ji} \] \[\Huge [A]'=[A]\]

OpenStudy (anonymous):

Good! So we are given 5 zeros, 2 negative ones, and 2 ones And we are asked to arrange this set of numbers into symmetric matices?

OpenStudy (dls):

Absolutely :)

OpenStudy (anonymous):

And we are looking for the number of symmetric matrices, right?

OpenStudy (anonymous):

Ok. The first thing to do is determine the size of the matrix.

OpenStudy (anonymous):

@DLS Like you said, for a symmetric matrix, we need matrix A to equal the transpose matrix A(T)

OpenStudy (anonymous):

@DLS if this is the case, what can you tell me about the size of the matrix?

OpenStudy (dls):

Yes

OpenStudy (anonymous):

hat can you tell me about the size of the matrix?

OpenStudy (anonymous):

what*

OpenStudy (dls):

3x3

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

Ok, let's set up a generic 3x3 matrix and it's transpose and see if we can go from there.

OpenStudy (anonymous):